# revtex4-1Repair the float

### v1

@@ I Introduction
Majorana zero modes (MZMs) emerge in topological superconductors as charge neutral quasiparticles [1, 2, 3, 4] and have attracted an enormous interest due to their potential for quantum computing applications [5, 6, 7, 8]. In one dimension, topological superconductivity with MZMs has been shown to appear in the so-called Kitaev chain [9], which consists of spin-polarized fermions with $p$-wave superconductivity. Although this type of superconductivity is scarce in nature [10], it was predicted to occur by combining conventional ingredients such as spin-singlet $s$-wave superconductivity, spin-orbit coupling, and a magnetic field [11, 12]. The simplicity of this proposal has motivated several studies aiming to detect MZMs but, do date, there is no consensus on whether MZMs have been observed or not [13].
Part of the challenges is believed to be due to the complex experimental setups [1, 2], which inevitably enable the presence of other phenomena that obscures Majorana physics [13]. To mitigate some of the issues, Kitaev chains with few sites are now being pursued [14, 15, 16, 17], which offers a bottom-up engineering of a Kitaev chain. In this case, MZMs emerge at single points (sweet spots) in the parameter space and do not exhibit any topological protection [18], properties that made them to be coined poor man’s Majorana modes (PMMMs). Interestingly, these PMMMs appear as charge neutral quasiparticles, having zero energy and being spatially nonlocal [19, 20], in the same way as MZMs in topological superconductors [4]. A unique consequence of the charge neutrality of MZMs is that it originates a divergent odd-frequency pairing that is intimately tied to their topology [21, 22, 23, 4], revealing that the superconducting pairing with MZMs is highly unusual, see also Refs. [24, 25, 26, 27, 28, 29, 30]. Since PMMMs exhibit a charge neutrality similar to MZMs but lack of topological protection [18, 19, 20], it is natural to wonder what is the nature of the emergent superconducting pairing in the presence of PMMMs.
@@ II Pair symmetries in few-site Kitaev chains
We begin by providing a general symmetry classification of the superconducting pair correlations allowed by all the quantum numbers in few-site Kitaev chains. To characterize the superconducting pair correlations, we employ the anomalous Green’s function defined as ${\mathcal{F}}^{nm}_{\alpha\beta}(t,t^{\prime})=\langle\mathcal{T}c_{\alpha n}(t)c_{\beta m}(t^{\prime})\rangle$ where $\mathcal{T}$ is the time ordering operator, $c_{n\alpha}(t)$ annihilates an electronic state in site $\alpha$ of superconductor $n$ at time $t$ [31, 32]. The anomalous Green’s function ${\mathcal{F}}^{nm}_{\alpha\beta}(t,t^{\prime})$ is also known as pair amplitude or pair correlation and its formation is fully tied to the quantum numbers of the two paired electrons, namely, $(n,m)$, $(\alpha,\beta)$, and time coordinates ($t,t^{\prime}$). For instance, the spin configuration of the Kitaev chain already dictates the spin symmetry of ${\mathcal{F}}^{nm}_{\alpha\beta}$ since we do not assume any active spin field: the Kitaev chain consists of spin-polarized fermions [9], which implies that all the fermions have the same spin and that the pair amplitude has a spin-triplet symmetry [24, 25, 26, 27, 28, 29, 30]; see also Ref. [33].
@@
The symmetries with respect to the other quantum numbers $(n,m)$, $(\alpha,\beta)$, and time coordinates ($t,t^{\prime}$) are not arbitrary but must respect the fermionic nature of the composed electrons. Since the pair amplitude behaves as a two-electron wavefunction, it must fulfill the antisymmetry condition such that ${\mathcal{F}}^{nm}_{\alpha\beta}(t,t^{\prime})=-{\mathcal{F}}^{mn}_{\beta\alpha}(t^{\prime},t)$ under the total exchange of quantum numbers plus exchange of time coordinates [21, 22, 34]. Similarly, in frequency domain, the antisymmetry condition implies ${F}^{nm}_{\alpha\beta}(\omega)=-{F}^{mn}_{\beta\alpha}(-\omega)$, where ${F}^{nm}_{\alpha\beta}(\omega)$ is the Fourier transform of ${\mathcal{F}}^{nm}_{\alpha\beta}(t,t^{\prime})$. Thus, the allowed pair symmetries must be antisymmetric under a total exchange of quantum numbers. Taking this into account, under the individual exchange of frequency ($\omega$), site indices $(\alpha,\beta)$, and superconductor indices $(n,m)$, the pair amplitude can be either an even (E) or odd (O) function. As pointed out in the previous paragraph, the Kitaev chain involves spin-polarized fermions which leaves the spin symmetry to be triplet (T). We find that there are four distinct spin triplet pair symmetries that obey the antisymmetry condition, see Tab.  1. It can be seen that the superconductor index $n$ (sup. index), as well as the dot index $\alpha$, play a role that is similar to the band index in multiband superconductor [34]. By focusing on the pair amplitudes inside a given superconductor, the sup. index restricts its symmetry to be even and only two pair symmetry classes are allowed: even-frequency, spin-triplet, odd under dot index, even under sup. index or ETOE; odd-frequency, spin-triplet, even under dot index, even under sup. index or OTEE. Thus, few-site Kitaev chains are expected to host distinct types of pair symmetry classes.
@@ III Emergent pair correlations in two-site Kitaev chains
where $\varepsilon_{\alpha}$ is the onsite energy of dot $\alpha=L/R$, $t$ is the hopping amplitude, and $\Delta$ the $p$-wave pair potential. Moreover, $\eta_{\pm}=(\eta_{0}\pm\eta_{z})/2$, with $\eta_{i}$ the $i$-th Pauli matrix in the site subspace, while $\tau_{i}$ the Pauli matrix in Nambu subspace. Notably, the minimal Kitaev chain given by Eq. 1 at the sweet spot $\varepsilon_{\rm L,R}=0$ and $\Delta=t$ hosts a pair of MZMs, with their wavefunctions fully located at the left and right sites but without being topologically protected, namely, Eq. 1 hosts a pair of PMMMs [18]. Moreover, it is worth noting that two-site Kitaev chain also holds experimental relevance as it has been recently realized in superconductor-semiconductor hybrids [14], see Fig. 1(a). Most of the properties of PMMMs have been shown to be similar to those of MZMs in topological superconductors. However, the nature of the induce superconducting pairing is still unknown. In particular, being the PMMMs of Majorana origin and having an intrinsic charge neutrality poses the question about the type of induced superconducting pairing.
@@ III.1 Majorana signatures in the spectral function
Before addressing the pair correlations, it is worth discussing the signatures of the PMMMs in the spectral function, obtained from the normal Green’s functions. By using Eq. (1) and Eq. (2), we find the normal Green’s function components given by
@@ III.2 Emergent pair amplitudes
Locally, the onsite pair amplitudes $F_{\rm LL(RR)}$ are evidently even under the exchange of the site index $\alpha=L,R$ and odd functions of frequency $\omega$; note that $D(\omega)=D(-\omega)$ is an even function of $\omega$ and $F_{\rm LL(RR)}$ exhibit opposite sign. Thus, the local pair amplitudes exhibit an odd-frequency, spin-triplet, even-site, even-sup. symmetry, which corresponds to the symmetry class OTEE of Tab. 1. This pair symmetry class is the only type of local emergent superconducting pairing emerging in these type of systems; note that the two-site Kitaev chain does not have a local superconducting pair potential, revealing that the OTEE class is indeed an emergent superconducting pairing. When it comes to the nonlocal pair amplitudes $F^{\pm}_{\rm LR}$, the expressions given in Eqs. (4) are already symmetrized with respect to the site index: $F^{+}_{\rm LR}$ is even under the exchange of L and R, while $F^{-}_{\rm LR}$ is odd. Moreover, Eqs. (4) already reveal that $F^{+}_{\rm LR}$ and is an odd function of $\omega$ while $F^{-}_{\rm LR}$ is an even function of $\omega$. With these considerations, together with the spin-triplet and even-sup. index symmetries, it is already clear that $F^{+}_{\rm LR}$ and $F^{-}_{\rm LR}$ correspond to the OTEE and ETOE symmetry classes of Tab. 1. Thus, nonlocally, the two-site Kitaev chain hosts two types of superconducting correlations, with the OTEE being entirely induced while the ETOE pairing tied to the parent $p$-wave pair potential.
@@ III.3 Pair amplitudes at the sweet spot $\Delta=t$
which at low frequencies can be approximated by $F_{\rm LL/RR}(\omega)\approx\mp[1/(2\omega)]\mp[\omega/(8\Delta)]$. It is thus evident the emergence of local pair amplitudes with a divergent profile near zero frequency [Fig. 3(a,d)]. Since the local pair amplitudes correspond to the OTEE pair symmetry class, their odd-frequency symmetry is fully determined by having a divergent frequency dependence. As noted in the introduction, having divergent odd-frequency pair amplitudes was shown to be a unique property of topological superconductors with MZMs, associated topology and revealing the Majorana nonlocality [24, 25, 26, 27, 28, 29, 30]; see also Refs. [21, 22, 23, 4]. In the present case, however, our minimal two-site Kitaev chain in the sweet spot hosts PMMMs without any topological properties but, surprisingly, we find divergent odd-frequency pairing. This occurs because even in the two-site Kitaev chain, it is possible to realize a pair of fully nonlocal zero-energy PMMMs in the same way as MZMs.
@@
Thus, at vanishing one onsite energy (e.g., $\varepsilon_{\rm R}=0$), the OTEE class $|F^{+}_{\rm LR}|$ exhibits large values at zero frequency as a function of $\varepsilon_{\rm L}$ but vanishes at $\varepsilon_{\rm L}=0$, see Fig. 4(a) and the green arrows; see also gray curve in Fig. 4(c). The large zero frequency values of $|F^{+}_{\rm LR}|$ remain robust under variations of $\varepsilon_{\rm L}\neq 0$, provided $\varepsilon_{\rm R}=0$; for $\varepsilon_{\rm R}\neq 0$, $|F^{+}_{\rm LR}|$ gets large values only near $\omega=0$, see also Fig. 4(b). Thus, when one of the onsite energies is finite and the other vanishes, $|F^{+}_{\rm LR}|$ develops a divergent profile near zero frequency, see red and green curves in Fig. 4(c) and also gray curve in Fig. 4(d). This divergent profile is similar to what we found for the local pair amplitude discussed in previous subsection. For finite onsite energies, however, no divergent profile is obtained: $|F^{+}_{\rm LR}|$ is peaked at the frequencies of the energy levels occurring away from zero frequency and vanish at $\varepsilon_{\rm R}=\varepsilon_{\rm L}$, see green arrows in Fig. 4(b) and red curve in Fig. 4(d). Away from this vanishing value but at distinct onsite energies, we find $|F^{+}_{\rm LR}|$ to depend linearly on $\omega$ when approaching zero frequency, see green curve in Fig. 4(d) and third expression in Eqs. (5).
@@
We have therefore obtained that two-site Kitaev chains in the sweet spot $\Delta=t$ exhibit local OTEE pair amplitudes with a divergent odd-frequency dependence around zero frequency, provided one or both onsite energies vanish. Moreover, the nonlocal OTEE pair symmetry class also exhibits a divergent odd-frequency dependence around zero frequency, as long as either of the onsite energies vanishes and $\varepsilon_{\rm L}\neq\varepsilon_{\rm R}$. Thus, the divergent odd-frequency profile of the emergent local and nonlocal pair correlations constitute a characteristic of the unconventional superconducting state with PMMMs.

### v2

@@ I Introduction
Majorana zero modes (MZMs) emerge in topological superconductors as charge neutral quasiparticles [1, 2, 3, 4] and have attracted an enormous interest due to their potential for quantum computing applications [5, 6, 7, 8]. In one dimension, topological superconductivity with MZMs has been shown to appear in the so-called Kitaev chain [9], which consists of spin-polarized fermions with $p$-wave superconductivity. Although this type of superconductivity is scarce in nature [10], it was predicted to occur by combining conventional ingredients such as spin-singlet $s$-wave superconductivity, spin-orbit coupling, and a magnetic field [11, 12]. The simplicity of this proposal has motivated several studies aiming to detect MZMs but, to date, there is no consensus on whether MZMs have been observed or not [13].
Part of the challenges is believed to be due to the complex experimental setups [1, 2], which inevitably enable the presence of other phenomena that obscures Majorana physics [13]. To mitigate some of the issues, Kitaev chains with few sites are now being pursued [14, 15, 16, 17], thus offering a bottom-up engineering approach. In this case, MZMs emerge at fine-tuned single points in the parameter space, commonly referred to as sweet spots, but do not exhibit any topological protection [18]. For this reason, such MZMs were coined as poor man’s Majorana modes (PMMMs); see also Ref. [19]. Interestingly, these PMMMs appear as charge neutral quasiparticles, having zero energy and being spatially nonlocal [20, 21], in the same way as MZMs in topological superconductors [4]. A unique consequence of the charge neutrality of MZMs is that it originates a divergent odd-frequency pairing that is intimately tied to their topology [22, 23, 24, 4], revealing that the superconducting pairing with MZMs is highly unusual, see also Refs. [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35].
Under general conditions, the superconducting pairing can have even- or odd-frequency symmetries, where the paired electrons forming Cooper pairs have a pair amplitude that is even or odd in their relative time, or frequency, see e. g.  Refs. [36, 23]. Identifying the pair symmetries can therefore help understanding the type of emergent superconducting pairing. Since PMMMs exhibit a charge neutrality similar to MZMs but lack of topological protection [18, 20, 21], it is natural to wonder what is the nature of the emergent superconducting pairing in the presence of PMMMs.
@@ II Pair symmetries in few-site Kitaev chains
We begin by providing a general symmetry classification of the superconducting pair correlations allowed by all the quantum numbers in few-site Kitaev chains. To characterize the superconducting pair correlations, we employ the anomalous Green’s function defined as ${\mathcal{F}}^{nm}_{\alpha\beta}(t,t^{\prime})=\langle\mathcal{T}c_{\alpha n}(t)c_{\beta m}(t^{\prime})\rangle$ where $\mathcal{T}$ is the time ordering operator, $c_{n\alpha}(t)$ annihilates an electronic state in site $\alpha$ of superconductor $n$ at time $t$ [37, 38]. The anomalous Green’s function ${\mathcal{F}}^{nm}_{\alpha\beta}(t,t^{\prime})$ is also known as pair amplitude or pair correlation and its formation is fully tied to the quantum numbers of the two paired electrons, namely, $(n,m)$, $(\alpha,\beta)$, and time coordinates ($t,t^{\prime}$). For instance, the spin configuration of the Kitaev chain already dictates the spin symmetry of ${\mathcal{F}}^{nm}_{\alpha\beta}$ since we do not assume any active spin field: the Kitaev chain consists of spin-polarized fermions [9], which implies that all the fermions have the same spin and that the pair amplitude has a spin-triplet symmetry [25, 26, 27, 28, 30, 31, 32, 33]; see also Ref. [39].
@@
The symmetries with respect to the other quantum numbers $(n,m)$, $(\alpha,\beta)$, and time coordinates ($t,t^{\prime}$) are not arbitrary but must respect the fermionic nature of the composed electrons. Since the pair amplitude behaves as a two-electron wavefunction, it must fulfill the antisymmetry condition such that ${\mathcal{F}}^{nm}_{\alpha\beta}(t,t^{\prime})=-{\mathcal{F}}^{mn}_{\beta\alpha}(t^{\prime},t)$ under the total exchange of quantum numbers plus exchange of time coordinates [22, 23, 40]. Similarly, in frequency domain, the antisymmetry condition implies ${F}^{nm}_{\alpha\beta}(\omega)=-{F}^{mn}_{\beta\alpha}(-\omega)$, where ${F}^{nm}_{\alpha\beta}(\omega)$ is the Fourier transform of ${\mathcal{F}}^{nm}_{\alpha\beta}(t,t^{\prime})$. Thus, the allowed pair symmetries must be antisymmetric under a total exchange of quantum numbers. Taking this into account, under the individual exchange of frequency ($\omega$), site indices $(\alpha,\beta)$, and superconductor indices $(n,m)$, the pair amplitude can be either an even (E) or odd (O) function. As pointed out in the previous paragraph, the Kitaev chain involves spin-polarized fermions which leaves the spin symmetry to be triplet (T). We find that there are four distinct spin triplet pair symmetries that obey the antisymmetry condition, see Tab.  1. It can be seen that the superconductor index $n$ (sup. index), as well as the dot index $\alpha$, play a role that is similar to the band index in multiband superconductors [41, 42, 43, 44, 40] and sideband index in Floquet superconductors [45, 46], see also Ref.  [^1] It is worth noting that the role of the dot index for broadening the emerging superconducting symmetries was studied before, see e.g. Refs. [54, 52, 53]. However, none of these previous works addressed few-site Kitaev chains.. The sup. index $n$ becomes active when multiple few-site Kitaev chains are coupled, e.g., in Josephson junctions [48], which are expected to be useful for realizing superconducting circuits. However, by focusing on the pair amplitudes inside a given superconductor, the sup. index restricts its symmetry to be even and only two pair symmetry classes are allowed: even-frequency, spin-triplet, odd under dot index, even under sup. index or ETOE; odd-frequency, spin-triplet, even under dot index, even under sup. index or OTEE. While in general all the allowed pair symmetries can appear as a linear combination, and can therefore coexist, the dominance of a particular symmetry characterizes the type of emergent superconducting pairing. Therefore, under general circumstances, few-site Kitaev chains are expected to host distinct types of pair symmetry classes.
@@ III Emergent pair correlations in two-site Kitaev chains
where $\varepsilon_{\alpha}$ is the onsite energy of dot $\alpha=L/R$, $t$ is the hopping amplitude, and $\Delta$ the $p$-wave pair potential. Moreover, $\eta_{\pm}=(\eta_{0}\pm\eta_{z})/2$, with $\eta_{i}$ the $i$-th Pauli matrix in the site subspace, while $\tau_{i}$ the Pauli matrix in Nambu subspace. Notably, the minimal Kitaev chain given by Eq. (1) at the sweet spot $\varepsilon_{\rm L,R}=0$ and $\Delta=t$ hosts a pair of MZMs, with their wavefunctions fully located at the left and right sites but without being topologically protected, namely, Eq. (1) hosts a pair of PMMMs [18]. Moreover, it is worth noting that two-site Kitaev chain also holds experimental relevance as it has been recently realized in superconductor-semiconductor hybrids [14], see Fig. 1(a). Most of the properties of PMMMs have been shown to be similar to those of MZMs in topological superconductors. However, the nature of the induce superconducting pairing is still unknown. In particular, being the PMMMs of Majorana origin and having an intrinsic charge neutrality poses the question about the type of induced superconducting pairing.
@@ III.1 Majorana signatures in the spectral function
Before addressing the pair correlations, it is worth discussing the formation of PMMMs. While this can be carried out in different ways, it is useful to explore the spectral function not only because it can be directly obtained from the normal Green’s functions obtained above but, importantly, also because it can be experimentally accessed via conductance measurements [49]. Since conductance is often measured in Majorana experiments [13, 14], the spectral signatures of PMMMs can help understanding their formation. By using Eq. (1) and Eq. (2), we find the normal Green’s function components given by
@@ III.2 Emergent pair amplitudes
Locally, the onsite pair amplitudes $F_{\rm LL(RR)}$ are evidently even under the exchange of the site index $\alpha=L,R$ and odd functions of frequency $\omega$; note that $D(\omega)=D(-\omega)$ is an even function of $\omega$ and $F_{\rm LL(RR)}$ exhibit opposite sign  [^2] The opposite signs of the local pair amplitudes ($F_{\rm LL}$ and $F_{\rm RR}$) have also been reported in the local pair amplitudes of finite topological superconductors with MZMs [4]; in our case, however, no topology is involved. Moreover, the Majorana polarization has also been shown to develop opposite signs at the edges of topological superconductors [55]. This suggests an intriguing relation between odd-frequency pairing and Majorana polarization.. Thus, the local pair amplitudes exhibit an odd-frequency, spin-triplet, even-site, even-sup. symmetry, which corresponds to the symmetry class OTEE of Tab. 1. This pair symmetry class is the only type of local emergent superconducting pairing emerging in these type of systems; note that the two-site Kitaev chain does not have a local superconducting pair potential, revealing that the OTEE class is indeed an emergent superconducting pairing. When it comes to the nonlocal pair amplitudes $F^{\pm}_{\rm LR}$, the expressions given in Eqs. (4) are already symmetrized with respect to the site index: $F^{+}_{\rm LR}$ is even under the exchange of L and R, while $F^{-}_{\rm LR}$ is odd. Moreover, Eqs. (4) already reveal that $F^{+}_{\rm LR}$ and is an odd function of $\omega$ while $F^{-}_{\rm LR}$ is an even function of $\omega$. With these considerations, together with the spin-triplet and even-sup. index symmetries, it is already clear that $F^{+}_{\rm LR}$ and $F^{-}_{\rm LR}$ correspond to the OTEE and ETOE symmetry classes of Tab. 1. Thus, nonlocally, the two-site Kitaev chain hosts two types of superconducting correlations, with the OTEE being entirely induced while the ETOE pairing tied to the parent $p$-wave pair potential.
@@ III.3 Pair amplitudes at the sweet spot $\Delta=t$
which at low frequencies can be approximated by $F_{\rm LL/RR}(\omega)\approx\mp[1/(2\omega)]\mp[\omega/(8\Delta)]$. It is thus evident the emergence of local pair amplitudes with a divergent profile near zero frequency [Fig. 3(a,d)]. Since the local pair amplitudes correspond to the OTEE pair symmetry class, their odd-frequency symmetry is fully determined by having a divergent frequency dependence. As noted in the introduction, having divergent odd-frequency pair amplitudes is a unique property of topological superconductors with MZMs, associated to topology and revealing the Majorana nonlocality [25, 26, 27, 28, 30, 31, 32, 33]; see also Refs. [22, 23, 24, 4]. In the present case, however, our minimal two-site Kitaev chain in the sweet spot hosts PMMMs without any topological properties but, surprisingly, we find divergent odd-frequency pairing. This occurs because even in the two-site Kitaev chain, it is possible to realize a pair of fully nonlocal zero-energy PMMMs in the same way as MZMs. The relation between Majorana nonlocality and divergent odd-frequency pairing can be further understood by noting that Majorana operators fulfil a self-conjugation property $\gamma_{i}^{\dagger}=\gamma_{i}$, which occurs when MZMs are fully nonlocal and charge neutral [4], namely, well separated from each other without any energy splitting and spatial overlap between them. Thus, taking into account $\gamma_{i}^{\dagger}=\gamma_{i}$, the anomalous pair correlation ($f$) of a Majorana operator can be written as $f\sim\langle\gamma_{i}\gamma_{i}^{\dagger}\rangle=\langle\gamma_{i}\gamma_{i}\rangle=1/\omega$. Thus, the Majorana nonlocality, inherited by both MZMs in topological superconductors and PMMMs in few-site Kitaev chains, naturally produce a divergent odd-frequency pairing as a strong signature of their unusual superconducting pairing.
@@
Thus, at vanishing one onsite energy (e.g., $\varepsilon_{\rm R}=0$), the OTEE class $|F^{+}_{\rm LR}|$ exhibits large values at zero frequency as a function of $\varepsilon_{\rm L}$ but vanishes at $\varepsilon_{\rm L}=0$, see Fig. 4(a) and the green arrows; see also gray curve in Fig. 4(c). The large zero frequency values of $|F^{+}_{\rm LR}|$ remain robust under variations of $\varepsilon_{\rm L}\neq 0$, provided $\varepsilon_{\rm R}=0$; for $\varepsilon_{\rm R}\neq 0$, $|F^{+}_{\rm LR}|$ gets large values only near $\omega=0$, see also Fig. 4(b). Thus, when one of the onsite energies is finite and the other vanishes, $|F^{+}_{\rm LR}|$ develops a divergent profile near zero frequency, see red and green curves in Fig. 4(c) and also gray curve in Fig. 4(d). This divergent profile is similar to what we found for the local pair amplitude discussed in previous subsection and can be understood to be a consequence of the Majorana nonlocality. This stems from the fact that, given distinct Majorana operators, obeying the self-conjugation Majorana property due to Majorana nonlocality, a pair correlation between two of them still yields a divergent odd-frequency nonlocal pairing, in the same way as it happens for the local pair amplitudes in Eqs. (6). Thus, having $F_{\rm LR}^{+}$ as a divergent odd-frequency nonlocal pairing in the presence of PMMMs can be interpreted as a measure of Majorana nonlocality. For finite onsite energies, however, no divergent profile is obtained: $|F^{+}_{\rm LR}|$ is peaked at the frequencies of the energy levels occurring away from zero frequency and vanish at $\varepsilon_{\rm R}=\varepsilon_{\rm L}$, see green arrows in Fig. 4(b) and red curve in Fig. 4(d). Away from this vanishing value but at distinct onsite energies, we find $|F^{+}_{\rm LR}|$ to depend linearly on $\omega$ when approaching zero frequency, see green curve in Fig. 4(d) and third expression in Eqs. (5).
@@
We have therefore obtained that two-site Kitaev chains in the sweet spot $\Delta=t$ exhibit local OTEE pair amplitudes with a divergent odd-frequency dependence around zero frequency, provided one or both onsite energies vanish. Moreover, the nonlocal OTEE pair symmetry class also exhibits a divergent odd-frequency dependence around zero frequency, as long as either of the onsite energies vanishes and $\varepsilon_{\rm L}\neq\varepsilon_{\rm R}$. Thus, the divergent odd-frequency profile of the emergent local and nonlocal pair correlations constitute a characteristic of the unconventional superconducting state with PMMMs. It is worth noting that local and nonlocal pair correlations are expected to play an important role in local and nonlocal transport [52, 53], which happens when attaching leads to the left and right sides of the two-site Kitaev chain studied here. In particular, Andreev reflection and crossed Andreev reflections can be directly determined from the local and nonlocal squared pair amplitudes [52, 53], respectively. In this regard, having local and nonlocal pair correlations with distinct functionalities, as we obtain here in the presence of PMMMs, is expected to induce distinct conductance signatures that would allow to identify the type of dominant superconducting pairing.